P=20-2Q demand function and P=10+6Q Supply function calculate consumers' and producer's surplus



1
Expert's answer
2022-01-18T15:12:19-0500

In equilibrium:

20 - 2Q = 10 + 6Q,

Q = 1.25 unit,

P = 20 - 2×1.25 = 17.5.

Consumer surplus is:

"CS = 0.5\u00d71.25\u00d7(20 - 17.5) = 1.5625."

"PS = 0.5\u00d71.25\u00d7(17.5 - 10) = 4.6875."